# Wigner Cat Phases: A finely tunable system for exploring the transition to quantum chaos

### v1

@@ 4 Numerical Experiments
Numerical investigation of how tuning mGOE gives rise to a deviation from limiting value $\mu=1.0$. At this limiting value ETH is valid in the corresponding representative quantum system’s dynamics. Decreasing $\mu$ parameter represents a more localized states, i.e., MBL. We use matrix order of $N=1000,500$ with ensemble sample size of $M=100$ for quantifying the uncertainties.
@@ 4.2 Spectral Nearest-Neigbour Spacings
Nearest-neighbour spacing for different $\mu$ tuning are shown in Figures 2a-2d for $\mu=0.54,0.70,0.86,0.98$ respectively. Deviation from Wigner-Dyson distribution is demonstrated at smaller $\mu$ values, with lower values showing heavy-tailed distribution. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error bars. We see that Wigner-Dyson distribution is recovered at small mixtures, i.e., higher $\mu$ values. Results supports the notion that MBL phases with heavy-tailed nearest-neighbour spacings are possible.
4.3 Transitions to quantum chaos: Adjecent Gap Ratios
We computed both mean values and distribution of adjacent gap ratios due to Oganesyan-Huse. An advantage of this measure is that, it doesn’t require unfolding the spectra as it isn’t sensitive to the unfolding procedure.
@@ 4.3 Transitions to quantum chaos: Adjecent Gap Ratios
Mean values of gap ratios inform us about the localisation of the quantum dynamics. Variation of degree of mixture $\mu$ describes the transition from quantum chaos to localisation. mGOE’s can generate such transition in a linear fashion, this is quantified in Figure 4.
@@ 5 Conclusion
Pedagogically easy to understand ensemble of random matrices is introduced for simulation of localization transition to quantum chaos. Contributions of such mixed ensemble are three fold: (1) advances the notion of matrix ensemble to arbitrarily shaped observable matrices, such as complex networks with varying size sub-components or quantum system as in this work, (2) provides physically plausable toy model that is easy to simulate the transition to quantum chaos with an arbitrarily precise fine tunable parameter for pedagogical and research purposes, and (3) we demonstrated that MBL phases with heavy-tailed nearest-neighbour spacings are possible that still yields mean adjacency-gap value of an integrable system, our results providing spectral data albeit simulated for a new MBL phase, i.e., heavy-tailed localized.
@@
We are grateful to Y. Süzen for her kind support for the quantum dynamics project. Especially for her encouragement in releasing the computational works as free-software in a Python package for educational use, as well as her recommendations on pedagogical presentation.
@@
Software toolkit [suzen25ley] and dataset is available.

### v2

@@ 4 Numerical Experiments
Numerical investigation of how tuning mGOE gives rise to a deviation from limiting value $\mu=1.0$. At this limiting value ETH is valid in the corresponding representative quantum system’s dynamics. Decreasing $\mu$ parameter represents a more localized states, i.e., MBL. We use matrix orders of $N=1000,500$ with ensemble sample size of $M=100$ for quantifying the uncertainties.
@@ 4.2 Spectral Nearest-Neigbour Spacings
Nearest-neighbour spacing for different $\mu$ tuning are shown in Figures 2a-2d for $\mu=0.54,0.70,0.86,0.98$ respectively. Deviation from Wigner-Dyson distribution is demonstrated at smaller $\mu$ values, with lower values showing heavy-tailed distribution. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error bars. We see that Wigner-Dyson distribution is recovered at small mixtures, i.e., higher $\mu$ values. Results supports the notion that MBL phases with heavy-tailed nearest-neighbour spacings are possible. We don’t observe full integrability even in small $\mu$ values based on spacing distributions, i.e., Poisson distribution is not observed. This indicates mGOE represents systems that can be tuned from heavy-tailed localisation, and localisation to quantum chaos: displaying range of localisation phases, Wigner Cat Phases.
4.3 Transitions to quantum chaos: Adjacent Gap Ratios
We computed both mean values and distribution of adjacent gap ratios due to Oganesyan-Huse. An advantage of this measure is that it doesn’t require unfolding the spectra as it isn’t sensitive to the unfolding procedure.
@@ 4.3 Transitions to quantum chaos: Adjacent Gap Ratios
Mean values of gap ratios inform us about the localisation of the quantum dynamics. Variation of degree of mixture $\mu$ describes the transition from quantum chaos to localisation. mGOE’s can generate such transition in a linear fashion: this is quantified in Figure 4.
@@ 5 Conclusion
Pedagogically easy to understand ensemble of random matrices is introduced for simulation of localization transition to quantum chaos. Contributions of such mixed ensemble are three fold: (1) advances the notion of matrix ensemble to arbitrarily shaped observable matrices, such as complex networks with varying size sub-components or disordered quantum system as in this work, (2) provides physically plausible single matrix ensemble that is easy to simulate the transition to quantum chaos with an arbitrarily precise fine tunable parameter for pedagogical and research purposes, and (3) we demonstrated that MBL phases with heavy-tailed nearest-neighbour spacings are possible that still yields mean adjacency-gap value of an integrable system hinting limitations of gap ratio statistics, our results providing spectral data albeit simulated for a new MBL phase, i.e., heavy-tailed localized.
@@
We are grateful to Y. Süzen for her kind support for the quantum dynamics project. Especially for her encouragement in releasing the computational works as free-software in a Python package Leymosun [60] for educational use, as well as her recommendations on pedagogical presentation. We also thank Dr. cand. Devanshu Shekhar from Indian Institute of Technology, Kharagpur for the helpful comments and discussion.
@@
Software toolkit Leymosun [60] is publicly available to generate the results.